[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Oct 15, 2016

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Oct 15 20:15:03 UTC 2016


HSU Shift Log for Sat Oct 15 2016 16:15:00 GMT-0400 (EDT)

Shift: H

Forecaster: J Beven

Technical Issues Summary:

ATCF issues: None.

N-AWIPS/AWIPS2 issues: OPC mentioned an observation from a drifting buoy  
near Nicole that appeared in their and in WPC's N-AWIPS system. This ob did  
not appear in our N-AWIPS system. Why?

Other Techincal issues: None.

Miscellaneous issues: MSG-3/METEOSAT-10 outage at about 1230 UTC. The  
latest news is that MET-10 data is flowing again, but it is using the  
header for MET-9.
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20161015/b4e137c0/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list